y=17-6.5(x+11) what is the slope model, what is the y-intercept, name another point on the line

Answer by checkley71(8403)

Y=17-6.5(X+11)
Y=17-6.5X-71.5
Y=-6.5X-54.5 THIS LINE HAS A SLOPE OF -6.5 & HAS A Y INTERCEPT OF (0,-54.5)
SETTING Y=0 WE GET:
0=-6.5X-54.5
-6.5X=54.5
X=54.5/-6.5
X=8.3846 OR (8.3846,0) FOR A SECOND POINT ON THIS LINE.